Sol Summary Report – January 30th


Crew 274 Sol Summary Report 01-31-2023

Sol: 1

Summary Title: Training Day – Extra-vehicular activities (EVA)and Martian Land Rovers

Author’s name: Sarah E. Guthrie “Ceres”, Commander

Mission Status: Possible weather impact for two requested EVAs on 1-31-2023, standing by for MCC approval.

Sol Activity Summary: Crew completed required EVA and rover training at Marble Ritual. EVA 1 departed habitat at 10:25am which was 25 minutes past the requested deployment time. Complications with EVA suits caused delays (see Anomalies in work and Operations Report). EVA 1 consisted of crew members: CDR – Guthrie (EVA 1 Lead), O’Hara, Pender (HSO) and DiBernardo on rovers Spirit and Perseverance. EVA 1 returned to habitat at 10:59am to meet the approved EVA request. EVA 2 deployed on time at 1:00pm with crew members Lojek (EVA 2 Lead), Loy, Pena and Hines on rovers Curiosity and Opportunity. EVA 2 returned to habitat at 1:55pm. The Solar Observatory was unveiled and inspected per MDRS Staff guidance and will be activated tomorrow. Crew experienced a small kitchen fire when the stovetop knob ceased to operate properly, MDRS Support guidance advised using pliers to turn the burner off. There was no damage to personnel or habitat. Additionally, Engineer Lojek evaluated the 500 gal water reservoir, it is estimated the crew has used approximately 45 gals of water in 24 hours. The crew is discussing water conservation efforts to extend the life of the reservoir to EOM.

Look Ahead Plan: Rain received this afternoon may impact EVA 3 and 4 requests due to muddy conditions. Should this occur, Crew 274 will shift mission schedule to habitat maintenance (cleaning), individual research projects, and mission planning for Sol 3 EVA requests.

Anomalies in work: EVA Suit #1 fan is inop. Engineer Lojek and Hab Specialist are investigating the issue. Oven light inop (noted in Operations Report).

Weather: Sunny, highs -0.5C, low -7C

Crew Physical Status: Optimal

EVA: EVA 1 and EVA 2 completed for training at Marble Ritual.

Reports to be filed: Sol Summary, Operations Report, Green Hab Report, Journalist Report, and Astronomer Report.

Support Requested: Food inventory complete, separate request to follow.

Copyright © The Mars Society. All rights reserved. | Main Site